OpenStudy (anonymous):

Using the starting salary data from able 3.3, and assuming an annual increase at the current cost of living (currently about 2%), calculate what you could expect to receive as a starting salary in your chosen field of study if you graduate three years from now. Bachelor’s Degrees Bioengineering / Biomedical$47,300

OpenStudy (phi):

to increase x by 2% you do x+ 2% * x 2% is 2/100 = 0.02 so this is the same as x+ 0.02x or (factor x out) x(1+0.02) = x*1.02 that is for 1 year to increase by 2% for 3 years do x* (1.02)^3 in this case, use 47,300 for x: 47300 * (1.02)^3

OpenStudy (anonymous):

thank you